Calculus I, Fall 2021 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework week 06

1. Section 3.7, problem 48 solution:

Figure 1: Section 3.7, problem 48 2. Section 3.7, problem 51(a) solution:

Figure 2: Section 3.7, problem 51(a)